OpenStudy (prominer10000):

Sorry that I'm posting lots of questions, I have 2 more after this but I'll say at least I actually provide how I would think I would solve the problem lol But anyway, same drill. I will post the expression on the comments below because I can't figure out how to put numbers with quotations in this question, can only seem to do it in the comments. What is the value of the expression?

OpenStudy (prominer10000):

\[2^{3} + 5^{2}\]

OpenStudy (prominer10000):

I'm thinking 2 x 2 x 2 which equals 6 and then 5 + 5 which equals ten. And then 6 + 10 which would equal 16.

OpenStudy (dumbcow):

No exponents imply multiplication 2^3 = 2*2*2 = 8 5^2 = 5*5 = 25

OpenStudy (mathmale):

2^3 is not equal to 6; it appears you are adding. No. Use 2 as a factor 3 times: 2*2*2.

OpenStudy (mathmale):

Use 5 as a factor twice: 5*5=?

OpenStudy (prominer10000):

So 2 x 2 x 2 = 8 and then 5 x 5 equals 25 so then do I just add these two together to get my answer? (33?)

OpenStudy (dumbcow):

yes thats what the "+" sign means

OpenStudy (prominer10000):

Ok thanks! :)